OpenStudy (anonymous):

Which statement is true about the product square root of 2(3square root of 2 + square root of 14)? A.) It is rational and equal to 12. B.) It is rational and equal to 18. C.) It is irrational and equal to 6 + 2square root of 18. D.) It is irrational and equal to 3 +square root of 18.

OpenStudy (solomonzelman):

\[\color{blue}{ 2(3\sqrt{2} +\sqrt{14} ) } \]\[\color{blue}{ 2(3\sqrt{2} +\sqrt{2}\sqrt{7} ) } \]\[\color{blue}{ 2\sqrt{2}~(3+\sqrt{7} ) } \] The choices to me seem abstruse, sorry.

OpenStudy (solomonzelman):

square root of 14, would be changed in any way.

OpenStudy (solomonzelman):

would NOT be, I meant.

OpenStudy (anonymous):

Its ok can you help with others

OpenStudy (solomonzelman):

Yeah, but you would have to open a new question for that.

OpenStudy (anonymous):

ok